OpenStudy (anonymous):

Find the equation of a line in standard form with a line passing through (2,1) and is parallel to the line 4x-3y=1

OpenStudy (anonymous):

first line is -3y = -4x +1 or y = (4/3) x -1/3 second line has same slope to be parallel y = (4/3) x + b find b from knowing this line goes through (2,1)